OpenStudy (anonymous):

For the function defined by: \[f(x) = \left[\begin{matrix}x^2 & x \le1 \\ 2x+1 & x>1\end{matrix}\right]\] evaluate f(0) and identify the function that will be used when x=0.

OpenStudy (anonymous):

Try using `\begin{cases}...\end{cases}`

OpenStudy (anonymous):

would f(0) be \[\left[\begin{matrix}0 & 0\le1 \\ 1 & 0\ge1\end{matrix}\right]\] ?

OpenStudy (anonymous):

no \(f(0)\) is one number

OpenStudy (anonymous):

is \(0\le 1\) or is \(0>1\) true?

OpenStudy (anonymous):

so what do i do if it says to evaluate f(0)?

OpenStudy (anonymous):

Since the top one's condition is true, it's expression represents \(f(0)\)

OpenStudy (anonymous):

answer my question above, then we will know

OpenStudy (anonymous):

@satellite73 i dont know exactly what you mean, but its in matrix form and told me to evaluate it by f(0)

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x^2 & \text{if} & x \leq 0 \\ 2x+1& \text{if} & x>0 \end{array} \right. \]

OpenStudy (anonymous):

this is what the question is!

OpenStudy (anonymous):

they omitted the word "if" which means, if the input is less than or equal to zero, use the top formula for example \[f(-5)=(-5)^2=25\] since \(-5<0\)

OpenStudy (anonymous):

whereas if the input is larger than zero, use the bottom formula for example \[f(10)=2\times 10+1=21\] since \(10>0\)

OpenStudy (anonymous):

since \(0\) is less than one, use the top formula

OpenStudy (anonymous):

i.e. use \(f(0)=0^2=0\)

OpenStudy (anonymous):

f(0) = 0^2=0, 0<=1, 2(0)+1=1, 0>1 would be the answer?

OpenStudy (anonymous):

no no the answer is one number

OpenStudy (anonymous):

the stuff to the right tells you what equation to use depending on what number you have

OpenStudy (anonymous):

so would the answer be just 0?

OpenStudy (anonymous):

yes, just zero

OpenStudy (anonymous):

i think you are confused by what this means it is called a piecewise function it is defined in different ways depending on the value of the input

OpenStudy (anonymous):

so what does it mean when it says to identify the function that will be used when x = 0?

OpenStudy (anonymous):

it means out of the two formulas \(x^2\) or \(2x+1\) which would you use

OpenStudy (anonymous):

since \(0<1\) you use the top one \(x^2\) and not the bottom one \(2x+1\)

OpenStudy (anonymous):

but if you wanted \(f(5)\) you would use the bottom one since \(5>1\)

OpenStudy (anonymous):

i.e. \(f(5)=2\times 5+1=11\) you wouldn't square it, since 5 is larger than 1

OpenStudy (anonymous):

i just dont understand where you're getting the numbers from and i dont know how to get the answers that i need D:

OpenStudy (anonymous):

i am making them up

OpenStudy (anonymous):

you need to learn how to read this to do the problem the stuff on the right tells you which formula to use

OpenStudy (anonymous):

lets just answer the quesion A) you would use the top expression \(x^2\) since \(0<1\) B) \(f(0)=0^2=0\)

OpenStudy (anonymous):

so if i use x^2, i get 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, how do i find the others now? :o

OpenStudy (anonymous):

what others?

OpenStudy (anonymous):

so to answer the question "identify the function that will be used when x=0" it would just be 0?

OpenStudy (anonymous):

because i have 2 questions. 1 that says "identify the function that will be used when x=0" & "find the value of this function when x=0" so what would the difference between the two answers be?

OpenStudy (anonymous):

i wrote the answers above exactly as A and B

OpenStudy (anonymous):

oh i see! im so sorry, thanks!!! :)

OpenStudy (anonymous):

now part II is saying "Graph f(x). Graph the first function completely."

OpenStudy (anonymous):

i know it is now f(0) but how do i graph this? @satellite73